OpenStudy (anonymous):

The ratio of the area of a smaller triangle to the area of a larger triangle is 1:4. If the area of the smaller triangle is 30, find the area of the larger triangle. Is the answer just 120?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It would be.. sqrt(30)=5.477 5.477*4=21.91 21.91*21.91=480

OpenStudy (anonymous):

But it's not a square, it's a triangle

OpenStudy (anonymous):

If it was a square its area would be sqrt(60)=7.745 7.745*7.745*(1/2)=30 5.477*5.477=30 So the answer is still correct.

OpenStudy (kinggeorge):

If it's a ratio of their areas, the area of the larger should be 120.

OpenStudy (anonymous):

So which one is it o-o

OpenStudy (anonymous):

Im sorry I should of explained this better. A=.5(lenght X width) 30=.5(7.745*7.745) The new triangle has sides that are 4 times as bigger. x=.5(30.98*30.98) x=480

OpenStudy (kinggeorge):

We're looking at a ratio of areas, not sides. Suppose you're first triangle has area of 30. Now we want to find a triangle that has an area x such that x/30=4/1. Solve this, you get that x=120.

OpenStudy (anonymous):

fight!fight!fight!fight!

OpenStudy (anonymous):

King is right -_-

OpenStudy (kinggeorge):

I've made worse mistakes than doing ratios wrong :P